# Analyzing GitHub Issues and Pull Requests in nf-core Pipelines: Insights into nf-core Pipeline Repositories

### v1

@@ 4.1. RQ1: What topics do users raise and discuss?
  • has-code presents a medium effect size ($\delta=0.50$), implying that the presence of code snippets differs meaningfully between open and closed issues and PRs.
@@
  • title-length ($\delta=0.28$) and has-URL ($\delta=0.37$) exhibit small effect sizes, showing modest differences.
@@
  • has-assignee ($\delta=0.09$), body-length ($\delta=0.001$), and num-contributors ($\delta=0.14$) have negligible effect sizes, indicating that although statistically significant, the differences in these features are not practically meaningful.
@@
Overall, the results highlight that labeling and the inclusion of code snippets are the most distinguishing features between open and closed issues and PRs, while differences in assignee presence, body length, and contributor count are less impactful.
RQ2 Summary RQ2 investigates how issues and PRs are managed and resolved, and which GitHub features influence their resolution. We find that 89.38% are closed, with nearly half resolved within three days. Feature analysis shows that repositories making active use of GitHub’s management mechanisms, labeling (used in 33.23% of cases) and assignment (17.13%), achieve higher closure rates. Statistical tests reveal that labeling (large effect, δ = 0.94) and including code snippets (medium effect, δ = 0.50) are the strongest predictors of successful resolution, while title length and URLs show smaller effects.
4.3. RQ3: Developers’ perceived difficulty of the identified topics.
@@
Motivation: Building on the topics identified from GitHub issues and PRs, our next step is to examine the difficulty of resolving challenges within each category. Not all issues carry the same weight: routine tasks like version updates or reporting and quality control visualization are relatively straightforward, whereas debugging containerized workflows, integrating genomic data, or maintaining pipeline quality often require deeper expertise, extensive coordination, and more time. Assessing perceived difficulty is important for several reasons. First, it highlights which categories impose the greatest cognitive and technical burden, guiding improvements in tooling, documentation, and contributor support. Second, it sheds light on the socio-technical complexity of maintaining the pipelines, where technical tasks must be balanced with collaborative processes. Moreover, it helps the community prioritize resources and streamline practices around the most demanding areas, thereby enhancing both the sustainability and usability of nf-core pipelines.
@@
Approach: We evaluate the difficulty of each topic using two complementary metrics: (1) the percentage of unresolved issues and PRs, and (2) the median time required for their resolution. This approach follows established practices in prior studies (Alam et al., 2025; Li et al., 2021; Scoccia et al., 2021). The details of these metrics are outlined below:
@@
1. The percentage of unresolved issues and PRs of a topic (% w/o solutions). The percentage of unresolved issues and PRs within a specific topic reflects how many are unresolved within that particular topic area. For each topic identified in RQ1, we calculate the proportion of issues and PRs that have not been closed. Closing typically indicates that a developer has addressed the problem or implemented the requested features, and thus considers it resolved. Consequently, topics with a higher proportion of open issues and PRs can be considered more difficult.
@@
2. The median time for issues and PRs to be resolved (Median time to resolve (hr)). We calculate the median resolution time (in hours) for issues and PRs based on the interval between their creation and closing times. This metric applies only to closed issues and PRs, as open ones lack a closing timestamp. A longer resolution time indicates greater difficulty in addressing the issue or implementing the requested changes.
@@
Results: Table 3 presents the proportion of unresolved issues and the median time required to close them across the 13 topics identified in RQ1. The results reveal substantial variation in both persistence and resolution speed, highlighting differing levels of technical and organizational complexity across topic categories.
The highest unresolved rate is observed in Tool Development and Repository Maintenance (20.28%), indicating that tasks related to introducing new tools, refactoring repositories, and maintaining cross-pipeline dependencies remain particularly challenging. Despite a relatively short median resolution time (36.83 h), the high proportion of unresolved cases suggests that while simpler maintenance tasks are closed quickly, complex tool-related efforts often stall, likely due to dependency conflicts, required domain expertise, or cross-pipeline dependencies. Similarly, Testing Pipelines and Managing CI Configurations (12.17%, 36.44 h) shows a notable difficulty level. Issues and PRs in this category frequently involve cloud-based execution, large datasets, or hardware constraints, making debugging and reproducibility time-consuming. Genome Data Integration and Testing (11.04%, 72.57 h) and Debugging Execution Failures (10.72%, 70.24 h) also exhibit long resolution times, reflecting the computational and data-specific challenges of running containerized workflows and managing diverse sequencing datasets.
@@
Overall, issues and PRs tied to integration, debugging, and infrastructure management remain the most persistent and time-intensive, whereas documentation, visualization, and versioning tasks are comparatively well-managed. These findings underscore the need for stronger automation, clearer contributor guidelines, and enhanced CI diagnostics to mitigate the high effort associated with complex technical tasks in nf-core pipeline maintenance.
RQ3 Summary Analysis of RQ3 shows clear variation in the difficulty of nf-core topics. The most demanding areas are Tool Development and Repository Maintenance (20.28%, 36.83 h), Testing Pipelines and Managing CI Configurations (12.17%, 36.44 h), and Genome Data Integration and Testing (11.04%, 72.57 h), reflecting the high technical and coordination effort required for integration, debugging, and infrastructure management. Similarly, Debugging Execution Failures (10.72%, 70.24 h) and Subworkflow Development and Integration (11.72%, 66.17 h) exhibit prolonged resolution times, underscoring their complexity. In contrast, routine tasks such as Reporting and Quality Control Visualization (9.44%, 29.37 h) are resolved more quickly.
@@ 5. Discussion
In this study, we analyze 25,173 issues and PRs from 125 nf-core pipelines, averaging 201.38 entries per repository. Among the pipelines, sarek (1,947), rnaseq (1,583), and eager (1,057) contain the largest number of issues and PRs. On average, repositories received 55.45 stars, 36.86 forks, and 151.55 watchers, with rnaseq emerging as the most popular (1,093 stars, 793 forks) and mitodetect attracting the most watchers (230). As shown in Figure 4, issue and PR activity have steadily increased, particularly after 2022, signaling a surge in community engagement. This growth underscores nf-core’s evolution into a collaborative ecosystem for pipeline development, where active reporting, maintenance, and contribution practices demonstrate the platform’s central role in promoting standardized, reusable, and community-driven bioinformatics pipelines.
Figure 4. Yearly distribution of issues and PRs
Our findings reveal key insights into the socio-technical dynamics of nf-core pipeline development. First, the topic taxonomy (RQ1) shows that discussions span the entire pipeline lifecycle, from scaffolding and data integration to debugging and documentation, highlighting nf-core’s dual role as both a software engineering and bioinformatics community. Unlike typical open-source projects, nf-core discussions emphasize reproducibility, container compatibility, and reference data management. Second, the analysis of issue and PR management (RQ2) shows high closure rates and quick resolutions but also reliance on informal or implicit coordination, with limited use of labels and assignees, suggesting the need for scalable governance as the ecosystem grows. Third, difficulty assessments (RQ3) identify tool development, CI configuration, and integration-heavy tasks as the most demanding, requiring stronger automation, diagnostics, and contributor guidance, whereas routine maintenance and documentation are handled efficiently, reflecting the maturity of nf-core’s community and infrastructure.
Collectively, these findings portray nf-core pipelines as a mature and sustainable collaborative ecosystem. Its evolution mirrors broader trends in open-source scientific infrastructure, where automation and community governance reduce maintenance burdens. Future work could extend this study by analyzing contributor trajectories, comparing nf-core pipelines to other OSS ecosystems, and exploring AI-based triage or summarization tools to further ease community workload.
@@ 6. Implications
For practitioners and maintainers: The study underscores the value of structured management practices. The strong association between labeling and resolution success suggests that systematic triaging can greatly enhance efficiency. Repositories should adopt clear labeling schemes, automated triage, and contributor assignment mechanisms. Given that tool development and CI configuration are especially challenging, maintainers should also offer better documentation, onboarding resources, and reusable templates to ease contributor participation.
For contributors: Detailed and well-structured reports are key to effective collaboration. Issues and PRs with longer descriptions and embedded code snippets show higher closure rates, indicating that contributors can enhance responsiveness by providing sufficient context, referencing related documentation, and adhering to standardized templates and reporting conventions.
For the broader community: The nf-core ecosystem demonstrates how a domain-specific community can achieve high responsiveness and sustainability by balancing automation with human expertise. However, reliance on informal coordination may present challenges as the community scales. Strengthening governance through clearer contribution guidelines, structured review processes, and community training will be essential for sustaining long-term growth.
For researchers: This work emphasizes the need for ongoing empirical studies of pipeline development. Our taxonomy of topics provides a foundation for comparative studies across domains (e.g., bioinformatics, AI, data science), while the identified difficulties highlight opportunities to develop automated tools for CI management, dependency resolution, and reproducibility assurance.
@@ 7. Threats to Validity
Internal validity: Our analysis relies on BERTopic to cluster GitHub issues and PRs. Different modeling choices, such as alternative clustering methods, preprocessing steps, or embedding models, could yield different topics. To ensure quality, we combine systematic preprocessing (e.g., lemmatization, stopword removal, improved vectorization) with manual validation. Three experienced researchers collaboratively review and label the generated topics, resolving disagreements through consensus. While this mitigates subjectivity, topic labeling involves some interpretive bias.
Construct Validity: Our study relies on GitHub, the most widely used platform for managing open-source issues and PRs for nf-core pipeline development. While GitHub provides rich insights, it may not capture all user challenges, particularly those discussed in external forums, slack channels, or mailing lists. We assess topic difficulty using unresolved issues/PRs and median resolution time. Although these indicators are informative, they may not fully capture all aspects of difficulty and thus pose a potential threat to construct validity. Other potential indicators, such as the number of comments, labels, review comments, and review cycles, could offer further insights. However, our metrics are consistent with those used in prior related studies (Alam et al., 2025; Li et al., 2021; Abdellatif et al., 2020).
@@
Conclusion Validity: Statistical associations, such as those between issue/PR characteristics and closure outcomes, should be interpreted cautiously. Although we use non-parametric tests and effect size measures to ensure robust comparisons, the presence of confounding variables (e.g., repository size, contributor experience) may still influence results. However, rigorous testing and transparent reporting support the reliability and validity of our conclusions.
@@ 8. Conclusion
This paper provides the first empirical analysis of 25,173 GitHub issues and pull requests across 125 active nf-core pipeline repositories to understand the challenges users face during the development and maintenance of these pipelines. Through BERTopic modeling, we identify 13 topics that span the entire pipeline lifecycle, from pipeline scaffolding and module updates to debugging, continuous integration, and quality control reporting. Our results show that 89.38% of issues and PRs are resolved, with half closed within three days, demonstrating an active and responsive community. Statistical analysis further reveals that labeling and inclusion of code snippets strongly influence resolution success, underscoring the importance of structured communication and transparent triage. We also find that tool development, CI configuration, and genome data integration are the most demanding areas, requiring deeper expertise and coordination. In contrast, documentation, visualization, and version management tasks are relatively more straightforward.
In the future, we intend to conduct surveys and interviews to uncover the reasons behind the lack of widespread adoption of issue and PR management features in nf-core pipeline repositories and to explore ways to increase their usage. Additionally, we will investigate the topics of discussion surrounding the development of pipeline using different scientific workflow management systems, such as Galaxy and SnakeMake, on different platforms like GitHub, Slack, external forums and mailing lists.
Replication Package
To support reproducibility, we have made the complete replication package for this study available on Zenodo (Random, 2025).

### v2

@@ 4.1. RQ1: What topics do users raise and discuss?
  • has-code presents a medium effect size ($\mathit{d}=0.50$), implying that the presence of code snippets differs meaningfully between open and closed issues and pull requests.
@@
  • title-length ($\mathit{d}=0.28$) and has-URL ($\mathit{d}=0.37$) exhibit small effect sizes, showing modest differences.
@@
  • has-assignee ($\mathit{d}=0.09$), body-length ($\mathit{d}=0.001$), and num-contributors ($\mathit{d}=0.14$) have negligible effect sizes, indicating that although statistically significant, the differences in these features are not practically meaningful.
@@
Overall, the results highlight that labeling and the inclusion of code snippets are the most distinguishing features between open and closed issues and pull requests, while differences in assignee presence, body length, and contributor count are less impactful.
[…]
4.3. RQ3: Perceived relevance of development activities to challenging topics.
@@
Motivation: Building on the topics, our next step is to examine the difficulty of resolving challenges within each category. Not all issues carry the same weight: routine tasks like version updates or reporting and quality control visualization are relatively straightforward, whereas debugging containerized workflows, integrating genomic data, or maintaining pipeline quality often require deeper expertise, extensive coordination, and more time. Assessing perceived difficulty is important for several reasons. First, it highlights which categories impose the greatest cognitive and technical burden, guiding improvements in tooling, documentation, and contributor support. Second, it sheds light on the socio-technical complexity of maintaining the pipelines, where technical tasks must be balanced with collaborative processes. Moreover, it helps the community prioritize resources and streamline practices around the most demanding areas, thereby enhancing both the sustainability and usability of nf-core pipelines.
@@
Approach: We evaluate the difficulty of each topic using two complementary metrics following established practices in prior studies (Alam et al., 2025; Li et al., 2021; Scoccia et al., 2021). The details of these metrics are outlined below:
@@
1. The percentage of unresolved issues and pull requests of a topic (% w/o solutions) reflects how many are unresolved within that particular topic area. For each topic identified in RQ1, we calculate the proportion of issues and pull requests that have not been closed. Closing typically indicates that a developer has addressed the problem or implemented the requested features and thus considers it resolved. Consequently, topics with a higher proportion of open issues and pull requests can be considered more difficult.
@@
2. The median time for issues and pull requests to be resolved (Median time to resolve (hr)). We calculate the median resolution time for issues and pull requests based on the interval between their creation and closing times. This metric applies only to closed issues and pull requests, as open ones lack a closing timestamp. A longer resolution time indicates greater difficulty in addressing the issue or implementing the requested changes.
@@
Results: Table 3 presents the proportion of unresolved issues and pull requests and the median time required to close them across the topics identified in RQ1. The results reveal substantial variation in both persistence and resolution speed, highlighting differing levels of complexity across topic categories.
The highest unresolved rate is observed in Tool Development and Repository Maintenance (20.28%), indicating that tasks related to introducing new tools, refactoring repositories, and maintaining cross-pipeline dependencies remain particularly challenging. Despite a relatively short median resolution time (36.83 h), the high proportion of unresolved cases suggests that while simpler maintenance tasks are closed quickly, complex tool-related efforts often stall, likely due to dependency conflicts, required domain expertise, or cross-pipeline dependencies. Similarly, Testing Pipelines and Managing CI Configurations (12.17%, 36.44 h) shows a notable difficulty level. Issues and pull requests in this category involve cloud-based execution, large datasets, or hardware constraints, making debugging and reproducibility time-consuming. Genome Data Integration and Testing (11.04%, 72.57 h) and Debugging Execution Failures (10.72%, 70.24 h) also exhibit long resolution times, reflecting the computational and data-specific challenges of running containerized workflows and managing diverse sequencing datasets.
@@
Overall, issues and pull requests tied to integration, debugging, and infrastructure management remain the most persistent and time-intensive, whereas documentation, visualization, and versioning tasks are comparatively well-managed. These findings underscore the need for stronger automation, clearer contributor guidelines, and enhanced CI diagnostics to mitigate the high effort associated with complex technical tasks in nf-core pipeline maintenance.
RQ3 Summary Analysis of RQ3 shows that the most complex areas are Tool Development and Repository Maintenance (20.28%, 36.83 h), Testing Pipelines and Managing CI Configurations (12.17%, 36.44 h), and Genome Data Integration and Testing (11.04%, 72.57 h), reflecting the high technical and coordination effort required for integration, debugging, and infrastructure management. Similarly, Debugging Execution Failures (10.72%, 70.24 h) and Subworkflow Development and Integration (11.72%, 66.17 h) exhibit prolonged resolution times, underscoring their complexity.
@@ 5. Discussion
In this study, we analyze 25,173 issues and pull requests from 125 nf-core pipelines, averaging 201.38 entries per repository. Among the pipelines, sarek (1,947), rnaseq (1,583), and eager (1,057) contain the largest number of issues and pull requests. On average, repositories received 55.45 stars, 36.86 forks, and 151.55 watchers, with rnaseq emerging as the most popular (1,093 stars, 793 forks) and mitodetect attracting the most watchers (230). As shown in Figure 4, issues and pull requests activity have steadily increased, particularly after 2022.
Figure 4. Yearly distribution of issues and pull requests
This growth underscores nf-core’s evolution into a collaborative ecosystem, where active maintenance, and contribution practices demonstrate the platform’s central role in promoting standardized, reusable, and community-driven bioinformatics pipelines.
Our findings reveal key insights into the nf-core pipeline development. First, the topic taxonomy (RQ1) shows that discussions span the entire pipeline lifecycle, highlighting nf-core’s dual role as both a software engineering and bioinformatics community. Unlike typical open-source projects, nf-core discussions emphasize reproducibility, container compatibility, and reference data management. Second, the analysis of issue and pull request management (RQ2) shows high closure rates and quick resolutions, but also reliance on informal or implicit coordination, suggesting the need for scalable governance as the ecosystem grows. Third, difficulty assessments (RQ3) identify tool development, CI configuration, and integration-heavy tasks as the most demanding, requiring stronger automation and contributor guidance. Collectively, these findings portray nf-core pipelines as a mature and sustainable collaborative ecosystem. Its evolution mirrors broader trends in open-source scientific infrastructure, where automation and community governance reduce maintenance burdens. Future work could extend this study by analyzing contributor trajectories, comparing nf-core pipelines to other OSS ecosystems, and exploring AI-based triage or summarization tools to further ease community workload.
@@ 6. Implications
For practitioners and maintainers: The study underscores the value of structured management practices. The strong association between labeling and resolution success suggests that systematic triaging can greatly enhance efficiency. Repositories should adopt clear labeling schemes, automated triage, and mechanisms for contributor assignment. Given that tool development and CI configuration are especially challenging, maintainers should also offer better documentation, onboarding resources, and reusable templates to ease contributor participation.
For contributors: Detailed and well-structured reports are key to effective collaboration. Issues and pull requests with longer descriptions and embedded code snippets show higher closure rates, indicating that contributors can enhance responsiveness by providing sufficient context and referencing related documentation.
For the broader community: The nf-core ecosystem shows how a domain-specific community can achieve high responsiveness and sustainability by balancing automation with human expertise. However, reliance on informal coordination may present challenges as the community scales. Strengthening governance through clearer contribution guidelines, structured review processes, and community training will be essential for sustaining long-term growth.
For researchers: Our taxonomy of topics provides a foundation for comparative studies across domains (e.g., bioinformatics, AI, data science), while the identified difficulties highlight opportunities to develop automated tools for CI management, dependency resolution, and reproducibility assurance.
@@ 7. Threats to Validity
Internal validity: Our analysis relies on BERTopic to cluster GitHub issues and pull requests. Different modeling choices, such as alternative clustering methods, preprocessing steps, or embedding models, could yield different topics. To ensure quality, we combine systematic preprocessing (e.g., lemmatization, stopword removal, improved vectorization) with manual validation. Three experienced researchers collaboratively review and label the generated topics, resolving disagreements through consensus. While this mitigates subjectivity, topic labeling involves some interpretive bias.
Construct Validity: Our study relies on data from GitHub, the most widely used platform for managing issues and pull requests in nf-core pipeline development. Although GitHub offers rich and structured insights, it may not fully capture all practitioner challenges, particularly those discussed in external sources. We assess topic difficulty using two indicators: unresolved issues and pull requests, and median resolution time. While these measures are informative, they may not fully capture all dimensions of difficulty, potentially threatening construct validity. Moreover, several confounding factors, including issue type, pipeline complexity and maturity, community engagement, and developer experience, may influence resolution outcomes. Additional indicators, such as the number of comments, labels, review comments, and review cycles, could provide further nuance. Nevertheless, the selected metrics are consistent with those employed in prior related studies (Alam et al., 2025; Li et al., 2021; Abdellatif et al., 2020).
@@
Conclusion Validity: Statistical associations should be interpreted cautiously. Although we use non-parametric tests and effect size measures to ensure robust comparisons, the presence of confounding variables (e.g., repository size, contributor experience) may still influence results. However, rigorous testing and transparent reporting support the reliability and validity of our conclusions.
@@ 8. Conclusion
This paper provides the first empirical analysis of 25,173 GitHub issues and pull requests across 125 active nf-core pipeline repositories to understand the challenges users face during the development and maintenance of these pipelines. Through BERTopic modeling, we identify 13 topics that span the entire pipeline lifecycle, from pipeline scaffolding and module updates to debugging, continuous integration, and quality control reporting. Our results show that 89.38% of issues and pull requests are resolved, with half closed within three days, demonstrating an active and responsive community. Statistical analysis further reveals that labeling and inclusion of code snippets strongly influence resolution success, underscoring the importance of structured communication and transparent triage. We also find that tool development, CI configuration, and genome data integration are the most demanding areas, requiring deeper expertise and coordination. In contrast, documentation, and version management tasks are relatively more straightforward.
In the future, we intend to conduct surveys and interviews to uncover the reasons behind the lack of widespread adoption of issues and pull requests management features in nf-core pipeline repositories and to explore ways to increase their usage. We also plan to investigate the topics of discussion surrounding the development of pipelines using different SWSs, such as Galaxy and Snakemake.
Acknowledgement
This research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR)